# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 2.2 $\mathcal{C}$-morphisms (or Lie-Bäcklund maps)
?proofname?.
The invariance condition (2) is equivalent to say that for any $\theta\in J^{\infty}(\pi)$ one has $\mathcal{B}_{*}(\left.D_{i}\right|_{\theta})=\sum_{j}a_{ij}\left.D_{j}\right|_{\mathcal{B}(\theta)}$, where $a_{ij}$ are some constants. Hence, for any function $f$ on $J^{\infty}(\pi^{\prime})$ one has that $D_{i}\left(\mathcal{B}^{*}(f)\right)(\theta)=\sum_{j}a_{ij}D_{j}(f)(\mathcal{B}(\theta))=\sum_{j}a_{ij}\mathcal{B}^{*}\left(D_{j}(f)\right)(\theta)$. Then, since this holds for any $\theta$ one readily gets (6).
Now, applying (6) to $f=x^{\prime}_{s}$ and using the fact that $\mathcal{B}^{*}\left(x^{\prime}_{s}\right)=\xi^{s}\left(x,u^{(k)}\right)$ and $\mathcal{B}^{*}\left(D_{j}^{\prime}\left(x^{\prime}_{s}\right)\right)=\mathcal{B}^{*}\delta_{js}=\delta_{js}$, one readily gets $D_{i}\xi^{s}=\alpha_{is}$ and hence (7) readily follows by (4) and (6). ∎
@@
Let $\mathcal{E}^{(\infty)}$ and $\mathcal{E}^{\prime}{}^{(\infty)}$ be the infinite prolongations of two formally integrable equations $\mathcal{E}\subset J^{k}(\pi)$ and $\mathcal{E}^{\prime}\subset J^{l}(\pi^{\prime})$, respectively. By a (regular) $\mathcal{C}$-morphism from $\mathcal{E}$ to $\mathcal{E}^{\prime}$ we mean a regular $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ such that $\mathcal{B}(\mathcal{E}^{(\infty)})\subseteq\mathcal{E}^{\prime}{}^{(\infty)}$.
@@
A regular $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ is a $\mathcal{C}$-morphism from =$\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}$ to =$\mathcal{E}^{\prime}=\{\mathbf{F^{\prime}}=\mathbf{0}\}$ if and only if
@@
?proofname?.
Since ()=$\mathcal{E}^{(\infty)}=\{D{}_{\sigma}\mathbf{F}=\mathbf{0}:|\sigma|\geq 0\}$ and =()$\mathcal{E}^{\prime}{}^{(\infty)}=\{D{}_{\rho}^{\prime}\mathbf{F}^{\prime}=\mathbf{0}:|\rho|\geq 0\}$, the condition $\mathcal{B}(\mathcal{E}^{(\infty)})\subseteq\mathcal{E}^{\prime}{}^{(\infty)}$ is equivalent to $\mathcal{B}^{*}\left(D^{\prime}_{\rho}\left(\mathbf{F}^{\prime}\right)\right)=0\mod{\mathcal{E}^{(\infty)}}$, for any multi-index $\rho$. Hence, in view of (7), $\mathcal{B}(\mathcal{E}^{(\infty)})\subseteq\mathcal{E}^{\prime}{}^{(\infty)}$ is equivalent to
@@
An example of regular $\mathcal{C}$-morphism between two equations is the Cole-Hopf transformation from the heat equation =$\mathcal{E}=\{u_{t}-u_{xx}=0\}$ to the Burgers equation =$\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xx}-2u^{\prime}u^{\prime}_{x}=0\}$, where $u=u(x,t)$ and $u^{\prime}=u^{\prime}(x,t)$. Indeed the Cole-Hopf transformation is defined by $u^{\prime}=u_{x}/u$, at the points where $u\neq 0$. In this case, by repeatedly applying (5), one can readily check that
@@
Another example of regular $\mathcal{C}$-morphism between two equations is the Miura transformation from the mKdV equation =$\mathcal{E}=\{u_{t}-u_{xxx}+6u^{2}u_{x}=0\}$ to the KdV equation =$\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xxx}-6u^{\prime}u^{\prime}_{x}=0\}$, where $u=u(x,t)$ and $u^{\prime}=u^{\prime}(x,t)$. For ease of comparison, here and in Example 14, it will be used the form of KdV adopted in [38] where the Miura transformation is defined by $u^{\prime}=u_{x}-u^{2}$. In this case, by repeatedly applying (5), one can readily check that
@@ 2.3.1 Symmetries of differential equations
Finite symmetries of a smooth distribution $\mathcal{D}$ on a manifold $N$ are diffeomorphisms $\psi:N\rightarrow N$ such that $\psi_{*}\mathcal{D}\subseteq\mathcal{D}$. Analogously, from the infinitesimal point of view, infinitesimal symmetries of a smooth distribution $\mathcal{D}$ on a manifold $N$ are smooth vector fields $Y$ on $N$ such that $L_{Y}\mathcal{D}\subseteq\mathcal{D}$. For instance, if $\mathcal{D}$ is generated by a system of vector fields, i.e., $\mathcal{D}=<X_{1},...,X_{n}>$, the symmetry condition is equivalent to $[Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}$, for any $i\in\{1,...,n\}$ and some smooth functions $\alpha_{i}^{s}$.
@@
Now, given a $k$-th order equation (or system) =$\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}\subset J^{k}(\pi)$, the classical finite symmetries of $\mathcal{E}$ are finite symmetries of the distribution $\mathcal{C}^{k}(\pi)$ which leave invariant the submanifold $\mathcal{E}$. Analogously, classical infinitesimal symmetries of $\mathcal{E}$ are vector fields on $J^{k}(\pi)$ which are infinitesimal symmetries of $\mathcal{C}^{k}(\pi)$ and are tangent to $\mathcal{E}$. Hence, by definition, the flow of an infinitesimal symmetry of $\mathcal{E}$ is a $1$-parameter local group of local diffeomorphisms which are finite symmetries of $\mathcal{E}$. A finite symmetry $\psi$ is called projectable if $\psi^{*}(C^{\infty}(M))\subset C^{\infty}(M)$. Analogously, an infinitesimal symmetry $X$ is called projectable if $X(C^{\infty}(M))\subset C^{\infty}(M)$.
@@
It can be seen (see [8] for details) that infinitesimal symmetries $X$ of $\mathcal{C}^{k}(\pi)$ are completely described by a generating function $\varphi=(\varphi^{j})$, defined as $\varphi^{j}:=X\,\lrcorner\,\omega^{j}$. This function is always of first order (i.e., only depend on $x_{i},u^{j},u_{i}^{j}$) and, whenever $m>1$, it is always of the form $\varphi^{a}=a_{i}(x_{i},u^{j})u_{i}^{a}+b^{a}(x_{i},u^{j})$. In particular, in canonical coordinates $\{x_{i},u_{\sigma}^{j}\}$, one has
$X=-\sum_{i}\frac{\partial\varphi^{a}}{\partial u_{i}^{a}}D_{i}^{(k)}+\sum_{|\sigma|=0}^{k-1}\,\sum_{j=1}^{m}D_{\sigma}^{(k)}(\varphi^{j})\partial_{u_{\sigma}^{j}},$ (10)
where $D_{\sigma}^{(k)}:=\left(D_{1}^{(k)}\right)^{\sigma_{1}}\circ...\circ\left(D_{n}^{(k)}\right)^{\sigma_{n}}$ and $a$ is any fixed integer in $\{1,...,m\}$.
@@
Thus, in view of (10), the higher order components of an infinitesimal symmetry can be obtained from lower order ones by means of a recurrence formula. In this sense one says that the infinitesimal symmetries of $\mathcal{C}^{k}(\pi)$ are prolongations of lower order vector fields. For instance, when $m>1$ any infinitesimal symmetry $X$ of $\mathcal{C}^{k}(\pi)$ is the prolongation of a vector field on $J^{0}(\pi)$; in such a case, $X$ is usually referred to as a point infinitesimal symmetry. On the other hand, when $m=1$ an infinitesimal symmetry $X$ of $\mathcal{C}^{k}(\pi)$ is in general the prolongation of a symmetry of $\mathcal{C}^{1}(\pi)$, that is not necessarily the prolongation of a vector field on $J^{0}(\pi)$. When $m=1$, $X$ is usually referred to as a contact infinitesimal symmetry whenever it is not a point symmetry.
@@
In practice, computing infinitesimal classical symmetries $X$ of an equation $\mathcal{E}\subset J^{k}(\pi)$ consists in the search of generating functions $\varphi=(\varphi^{j})=(X\,\lrcorner\,\omega^{j})$ such that $X$ is tangent to $\mathcal{E}$. This tangency condition leads to a linear system of PDEs for the function $\varphi$, which is usually overdetermined and hence can be algorithmically studied by taking into consideration the full set of compatibility conditions. The analysis of this kind of system is in general more feasible if one makes use of a symbolic manipulation package like those written for Maple.
@@
On the other hand, when $\mathcal{E}$ is formally integrable, a noteworthy extension of the notion of symmetry is possible. Indeed, one can say that a vector field $Z\in\mathcal{D}(\pi)$ is a symmetry of $\mathcal{E}^{(\infty)}$ if, and only if, $Z$ is a symmetry of $\mathcal{C}(\pi)$ which is tangent to $\mathcal{E}^{(\infty)}$. However, since $\mathcal{C}(\pi)$ is Frobenius, the Lie algebra $\mathcal{DC}(\pi)$ of vector fields inscribed in $\mathcal{C}(\pi)$ is an ideal of trivial (or characteristic) symmetries, that one wants preliminarily gauge out from symmetries of $\mathcal{C}(\pi)$. Thus, denoting by $\text{Sym}(\pi)$ the full Lie algebra of symmetries of $\mathcal{C}(\pi)$, one can define the Lie algebra of generalized (or higher) symmetries of $\mathcal{C}(\pi)$ as the quotient algebra $\text{sym}(\pi)=\text{Sym}(\pi)/\mathcal{DC}(\pi)$. Hence, since any field of $\mathcal{DC}(\pi)$ is already tangent to $\mathcal{E}^{(\infty)}$, all elements of a coset $\left[X\right]\in\text{sym}(\pi)$ (i.e., $\left[X\right]=X\,\mod{\mathcal{DC}(\pi)},$ with $X\in\text{Sym}(\pi)$) are tangent to $\mathcal{E}^{(\infty)}$, or neither are they. Thus, since it make sense to say that an element of $\text{sym}(\pi)$ is tangent or not to $\mathcal{E}^{(\infty)}$, one defines the Lie algebra of generalized (or higher) symmetries of $\mathcal{E}$ as the sub-algebra $\text{sym}(\mathcal{E})\subset\text{sym}(\pi)$ of generalized symmetries of $\mathcal{C}(\pi)$ that are tangent to $\mathcal{E}^{(\infty)}$.
@@
$X=\sum_{i}f_{i}D_{i}+\sum_{|\sigma|\geq 0}\sum_{j=1}^{m}D_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}},$ (11)
where $\varphi=(\varphi^{j})=(X\,\lrcorner\,\omega^{j})$ and $f_{i}$ are some differentiable functions. Thus, any higher symmetry $\left[X\right]$ can be identified with its vertical (or evolutionary) representative
@@
For any fixed $\varphi$, this is typically an overdetermined system for $\varphi$ that involves only a finite number of differential consequences of $\mathcal{E}$ and can be explicitly solved by considering its integrability conditions.
In particular, among the generalized symmetries of $\mathcal{E}$ one has the classical generalized symmetries, which are the vertical representatives of infinite prolongations of classical symmetries of $\mathcal{E}$.
For further details on the theory of classical and generalized symmetries, as well as for examples and explicit computations, see [8, 34].
2.3.2 Pseudosymmetries of differential equations
@@
In general, given a smooth distribution $\mathcal{D}$, a pseudosymmetry of $\mathcal{D}$ can be understood as a smooth vector field $Y$such that $L_{Y}\mathcal{D}\subseteq\,<Y>+\mathcal{D}$. For instance, for a distribution $\mathcal{D}=<X_{1},...,X_{n}>$, the pseudosymmetry condition is equivalent to say that for any $i\in\{1,...,n\}$ one has $[Y,X_{i}]=\alpha_{i}Y+\sum_{s}\beta_{i}^{s}X_{s}$, for some smooth functions $\alpha_{i}$ and $\beta_{i}^{s}$.
Sokolov introduced also the more general notion of an $r$-pseudosymmetry of $\mathcal{D}$, that can be seen as a system $\{Y_{1},...,Y_{r}\}$ of vector fields such that $L_{Y_{h}}\mathcal{D}\subseteq\,<Y_{1},...,Y_{r}>+\mathcal{D}$, for any $h=1,...,r$. Hence, for a distribution $\mathcal{D}=<X_{1},...,X_{n}>$, one has an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$ if for any $i\in\{1,...,n\}$ the $r\times 1$ matrix $\mathbb{Y}=\left[Y_{1},...,Y_{r}\right]^{T}$ is such that
@@
where $\left[\mathbb{Y},X_{i}\right]:=\left[\left[Y_{1},X_{i}\right],...,\left[Y_{r},X_{i}\right]\right]^{T}$ and $\mathbb{X}=\left[X_{1},...,X_{n}\right]^{T}$, whereas $\mathbb{U}_{i}$ and $\mathbb{V}_{i}$ are some $r\times r$ and $r\times n$ matrix-valued functions, respectively.
It is well known that the flow of an infinitesimal symmetry always transforms an integral manifolds of $\mathcal{D}$ to another (possibly different) integral manifolds. Also, in view of $[Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}$, it is readily seen that an infinitesimal symmetry of $\mathcal{D}$ sends an invariant (or first integral) $I$ of $\mathcal{D}$ to another invariant $Y(I)$. These properties are in general not true for pseudosymmetries. However, pseudosymmetries share with symmetries another important property described by the following
@@
Let $\left\{\xi^{1},...,\xi^{n},\nu^{1},...,\nu^{m}\right\}$ be invariants of a pseudosymmetry $Y$ (or, an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$) of $\mathcal{D}=<X_{1},...,X_{n}>$, such that
@@
$\left(\begin{array}[]{lll}\nu_{1}^{1}&\cdots&\nu_{1}^{m}\\ \vdots&&\vdots\\ \nu_{n}^{1}&\cdots&\nu_{n}^{m}\end{array}\right)=\left(\begin{array}[]{lll}X_{1}\left(\xi^{1}\right)&\cdots&X_{1}\left(\xi^{n}\right)\\ \vdots&&\vdots\\ X_{n}\left(\xi^{1}\right)&\cdots&X_{n}\left(\xi^{n}\right)\end{array}\right)^{-1}\left(\begin{array}[]{lll}X_{1}\nu^{1}&\cdots&X_{1}\nu^{m}\\ \vdots&&\vdots\\ X_{n}\nu^{1}&\cdots&X_{n}\nu^{m}\end{array}\right),$ (15)
@@
$\left(\begin{array}[]{lll}X_{1}\left(\xi^{1}\right)&\cdots&X_{1}\left(\xi^{n}\right)\\ \vdots&&\vdots\\ X_{n}\left(\xi^{1}\right)&\cdots&X_{n}\left(\xi^{n}\right)\end{array}\right)\left(\begin{array}[]{lll}\nu_{1}^{1}&\cdots&\nu_{1}^{m}\\ \vdots&&\vdots\\ \nu_{n}^{1}&\cdots&\nu_{n}^{m}\end{array}\right)=\left(\begin{array}[]{lll}X_{1}\nu^{1}&\cdots&X_{1}\nu^{m}\\ \vdots&&\vdots\\ X_{n}\nu^{1}&\cdots&X_{n}\nu^{m}\end{array}\right).$
@@
On the other hand, in view of $[Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}+\beta_{i}Y$ and the fact that $Y(\xi^{h})=Y(\nu^{j})=0$, last identity reduces to

### v5

@@ 2.2 $\mathcal{C}$-morphisms (or Lie-Bäcklund maps)
Proof.
The invariance condition (2) is equivalent to say that for any $\theta\in J^{\infty}(\pi)$ one has $\mathcal{B}_{*}(\left.D_{i}\right|_{\theta})=\sum_{j}a_{ij}\left.D_{j}\right|_{\mathcal{B}(\theta)}$, where $a_{ij}$ are some constants. Hence, for any function $f$ on $J^{\infty}(\pi^{\prime})$ one has that $D_{i}\left(\mathcal{B}^{*}(f)\right)(\theta)=\sum_{j}a_{ij}D_{j}f\,(\mathcal{B}(\theta))=\sum_{j}a_{ij}\mathcal{B}^{*}\left(D_{j}f\right)(\theta)$. Then, since this holds for any $\theta$ one readily gets (6).
Now, applying (6) to $f=x^{\prime}_{s}$ and using the fact that $\mathcal{B}^{*}\left(x^{\prime}_{s}\right)=\xi^{s}\left(x,u^{(k)}\right)$ and $\mathcal{B}^{*}\left(D_{j}^{\prime}\,x^{\prime}_{s}\right)=\mathcal{B}^{*}\delta_{js}=\delta_{js}$, one readily gets $D_{i}\xi^{s}=\alpha_{is}$ and hence (7) readily follows by (4) and (6). ∎
@@
[…]
@@
A regular $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ is a $\mathcal{C}$-morphism from $\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}$ to $\mathcal{E}^{\prime}=\{\mathbf{F^{\prime}}=\mathbf{0}\}$ if and only if
@@
Proof.
[…]
@@
An example of regular $\mathcal{C}$-morphism between two equations is the Cole-Hopf transformation from the heat equation $\mathcal{E}=\{u_{t}-u_{xx}=0\}$ to the Burgers equation $\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xx}-2u^{\prime}u^{\prime}_{x}=0\}$, where $u=u(x,t)$ and $u^{\prime}=u^{\prime}(x,t)$. Indeed the Cole-Hopf transformation is defined by $u^{\prime}=u_{x}/u$, at the points where $u\neq 0$. In this case, by repeatedly applying (5), one can readily check that
@@
Another example of regular $\mathcal{C}$-morphism between two equations is the Miura transformation from the mKdV equation $\mathcal{E}=\{u_{t}-u_{xxx}+6u^{2}u_{x}=0\}$ to the KdV equation $\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xxx}-6u^{\prime}u^{\prime}_{x}=0\}$, where $u=u(x,t)$ and $u^{\prime}=u^{\prime}(x,t)$. For ease of comparison, here and in Example 14, it will be used the form of KdV adopted in [41] where the Miura transformation is defined by $u^{\prime}=u_{x}-u^{2}$. In this case, by repeatedly applying (5), one can readily check that
@@ 2.3.1 Symmetries of differential equations
Finite symmetries of a smooth distribution $\mathcal{D}$ on a manifold $N$ are diffeomorphisms $\psi:N\rightarrow N$ such that $\psi_{*}\mathcal{D}\subseteq\mathcal{D}$. Analogously, from the infinitesimal point of view, infinitesimal symmetries of a smooth distribution $\mathcal{D}$ on a manifold $N$ are smooth vector fields $Y$ on $N$ such that $L_{Y}\mathcal{D}\subseteq\mathcal{D}$. Hence, the flow of an infinitesimal symmetry of $\mathcal{D}$ is a $1$-parameter local group of finite symmetries of $\mathcal{D}$. If $\mathcal{D}$ is generated by a system of vector fields, i.e., $\mathcal{D}=<X_{1},...,X_{n}>$, the symmetry condition is equivalent to $Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}$, for any $i\in\{1,...,n\}$ and some smooth functions $\alpha_{i}^{s}$.
@@
Now, given a $k$-th order equation (or system) $\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}\subset J^{k}(\pi)$, the classical finite symmetries of $\mathcal{E}$ are finite symmetries of the distribution $\mathcal{C}^{k}(\pi)$ which leave invariant the submanifold $\mathcal{E}$. Analogously, classical infinitesimal symmetries of $\mathcal{E}$ are vector fields on $J^{k}(\pi)$ which are infinitesimal symmetries of $\mathcal{C}^{k}(\pi)$ and are tangent to $\mathcal{E}$. A finite symmetry $\psi$ is called projectable if $\psi^{*}(C^{\infty}(M))\subseteq C^{\infty}(M)$. Analogously, an infinitesimal symmetry $X$ is called projectable if $X(C^{\infty}(M))\subseteq C^{\infty}(M)$.
@@
It can be seen (see [8] for details) that infinitesimal symmetries $X$ of $\mathcal{C}^{k}(\pi)$ are completely described by a generating function $\varphi=(\varphi^{j})$, defined as $\varphi^{j}:=X\,\lrcorner\,\omega_{0}^{j}$ (where $\omega_{0}^{j}=du^{j}-\sum_{i}u_{i}^{j}dx_{i}$). This function is always of first order (i.e., only depend on $x_{i},u^{j},u_{i}^{j}$) and, whenever $m>1$, it is always of the form $\varphi^{j}=\sum_{i}a_{i}(x,u)u_{i}^{j}+b^{j}(x,u)$. In particular, in canonical coordinates $\{x_{i},u_{\sigma}^{j}\}$, one has
$X=-\sum_{i}\frac{\partial\varphi^{s}}{\partial u_{i}^{s}}D_{i}^{(k)}+\sum_{|\sigma|=0}^{k-1}\,\sum_{j=1}^{m}D_{\sigma}^{(k)}\varphi^{j}\,\partial_{u_{\sigma}^{j}},$ (10)
where $D_{\sigma}^{(k)}:=\left(D_{1}^{(k)}\right)^{\sigma_{1}}\circ...\circ\left(D_{n}^{(k)}\right)^{\sigma_{n}}$ and $s$ is any fixed integer in $\{1,...,m\}$.
@@
Thus, in view of (10), the higher order components of an infinitesimal symmetry can be obtained from lower order ones by means of a recurrence formula. In this sense one says that the infinitesimal symmetries of $\mathcal{C}^{k}(\pi)$ are prolongations of lower order vector fields. For instance, when $m>1$ any infinitesimal symmetry $X$ of $\mathcal{C}^{k}(\pi)$ is the prolongation of a vector field on $J^{0}(\pi)$; in such a case, $X$ is usually referred to as a point infinitesimal symmetry. On the other hand, when $m=1$ an infinitesimal symmetry $X$ of $\mathcal{C}^{k}(\pi)$ is in general the prolongation of a symmetry of $\mathcal{C}^{1}(\pi)$, that is not necessarily the prolongation of a vector field on $J^{0}(\pi)$. In particular, when $m=1$, $X$ is usually referred to as a contact infinitesimal symmetry whenever it is not a point symmetry.
@@
Thus, the computation of infinitesimal classical symmetries $X$ of an equation $\mathcal{E}\subset J^{k}(\pi)$ reduces to the determination of functions $\varphi=(\varphi^{j})=(X\,\lrcorner\,\omega_{0}^{j})$ such that $X$ is tangent to $\mathcal{E}$. This condition provides typically an overdetermined linear system of PDEs for $\varphi$ that can be analyzed and explicitly solved by considering its integrability conditions. The analysis of such a system is generally more feasible if one uses some symbolic manipulation package of the type developed in a computer algebra system like Maple.
@@
On the other hand, when $\mathcal{E}$ is formally integrable, a noteworthy extension of the notion of symmetry is possible. Indeed, one can say that a vector field $X\in\mathcal{D}(\pi)$ is a generalised symmetry of $\mathcal{E}$ if, and only if, $X$ is a symmetry of $\mathcal{C}(\pi)$ which is tangent to $\mathcal{E}^{(\infty)}$. However, since $\mathcal{C}(\pi)$ is Frobenius, the Lie algebra $\mathcal{DC}(\pi)$ of vector fields inscribed in $\mathcal{C}(\pi)$ is an ideal of trivial (or characteristic) symmetries, that one wants preliminarily gauge out from symmetries of $\mathcal{C}(\pi)$. Thus, denoting by $\text{Sym}(\pi)$ the full Lie algebra of symmetries of $\mathcal{C}(\pi)$, one can define the Lie algebra of higher symmetries of $\mathcal{C}(\pi)$ as the quotient algebra $\text{sym}(\pi)=\text{Sym}(\pi)/\mathcal{DC}(\pi)$. Hence, since any field of $\mathcal{DC}(\pi)$ is already tangent to $\mathcal{E}^{(\infty)}$, all elements of a coset $\left[X\right]\in\text{sym}(\pi)$ (i.e., $\left[X\right]=X\,\mod{\mathcal{DC}(\pi)},$ with $X\in\text{Sym}(\pi)$) are tangent to $\mathcal{E}^{(\infty)}$, or neither are they. Thus, since it make sense to say that an element of $\text{sym}(\pi)$ is tangent or not to $\mathcal{E}^{(\infty)}$, one defines the Lie algebra of higher symmetries of $\mathcal{E}$ as the sub-algebra $\text{sym}(\mathcal{E})\subset\text{sym}(\pi)$ of higher symmetries of $\mathcal{C}(\pi)$ that are tangent to $\mathcal{E}^{(\infty)}$.
@@
$X=\sum_{i}a_{i}D_{i}+\sum_{|\sigma|\geq 0}\sum_{j=1}^{m}D_{\sigma}\varphi^{j}\,\partial_{u_{\sigma}^{j}},$ (11)
where $\varphi=(\varphi^{j})=(X\,\lrcorner\,\omega_{0}^{j})$ and $a_{i}$ are some differentiable functions. Thus, any higher symmetry $\left[X\right]$ can be identified with its vertical (or evolutionary) representative
@@
For any $\varphi$ of fixed order, this is typically an overdetermined linear system for $\varphi$ that involves only a finite number of differential consequences of $\mathcal{E}$ and can be explicitly solved by considering its integrability conditions. As with the symmetry condition for classical symmetries, the analysis of the linear system (13) is generally more feasible if one uses some symbolic manipulation package of the type developed in a computer algebra system like Maple.
For further details on the theory of classical and generalised symmetries, as well as for examples and explicit computations, see [8, 37].
2.3.2 Pseudosymmetries of differential equations and factorisation
@@
In general, given a smooth distribution $\mathcal{D}$, a pseudosymmetry of $\mathcal{D}$ can be understood as a smooth vector field YYsuch that $L_{Y}\mathcal{D}\subseteq\,<Y>+\mathcal{D}$. For instance, for a distribution $\mathcal{D}=<X_{1},...,X_{n}>$, the pseudosymmetry condition is equivalent to say that for any $i\in\{1,...,n\}$ one has $Y,X_{i}]=\alpha_{i}Y+\sum_{s}\beta_{i}^{s}X_{s}$, for some smooth functions $\alpha_{i}$ and $\beta_{i}^{s}$.
Sokolov introduced also the more general notion of an $r$-pseudosymmetry of $\mathcal{D}$, that can be seen as a system $\{Y_{1},...,Y_{r}\}$ of vector fields such that $L_{Y_{h}}\mathcal{D}\subseteq\,<Y_{1},...,Y_{r}>+\mathcal{D}$, for any h=1,,rh=1,...,r. Hence, for a distribution $\mathcal{D}=<X_{1},...,X_{n}>$, one has an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$ if for any $i\in\{1,...,n\}$ the $r\times 1$ matrix $\mathbb{Y}=\left[Y_{1},...,Y_{r}\right]^{T}$ is such that
@@
where $\left[\mathbb{Y},X_{i}\right]:=\left[\left[Y_{1},X_{i}\right],...,\left[Y_{r},X_{i}\right]\right]^{T}$ and $\mathbb{X}=\left[X_{1},...,X_{n}\right]^{T}$, whereas $\mathbb{U}_{i}$ and $\mathbb{V}_{i}$ are some $r\times r$ and $r\times n$ matrix-valued smooth functions, respectively.
It is well known that the flow of an infinitesimal symmetry always transforms an integral manifolds of $\mathcal{D}$ to another (possibly different) integral manifolds. Also, in view of $Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}$, it is readily seen that an infinitesimal symmetry of $\mathcal{D}$ sends an invariant (or first integral) $I$ of $\mathcal{D}$ to another invariant $Y(I)$. These properties are in general not true for pseudosymmetries. However, pseudosymmetries share with symmetries another important property described by the following
@@
Let $\left\{\xi^{1},...,\xi^{n},\nu^{1},...,\nu^{p}\right\}$ be invariants of a pseudosymmetry $Y$ (or, an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$) of $\mathcal{D}=<X_{1},...,X_{n}>$, such that
@@
$\left(\begin{array}[]{lll}\nu_{1}^{1}&\cdots&\nu_{1}^{p}\\ \vdots&&\vdots\\ \nu_{n}^{1}&\cdots&\nu_{n}^{p}\end{array}\right)=\left(\begin{array}[]{lll}X_{1}\left(\xi^{1}\right)&\cdots&X_{1}\left(\xi^{n}\right)\\ \vdots&&\vdots\\ X_{n}\left(\xi^{1}\right)&\cdots&X_{n}\left(\xi^{n}\right)\end{array}\right)^{-1}\left(\begin{array}[]{lll}X_{1}(\nu^{1})&\cdots&X_{1}(\nu^{p})\\ \vdots&&\vdots\\ X_{n}(\nu^{1})&\cdots&X_{n}(\nu^{p})\end{array}\right),$ (15)
@@
$\left(\begin{array}[]{lll}X_{1}\left(\xi^{1}\right)&\cdots&X_{1}\left(\xi^{n}\right)\\ \vdots&&\vdots\\ X_{n}\left(\xi^{1}\right)&\cdots&X_{n}\left(\xi^{n}\right)\end{array}\right)\left(\begin{array}[]{lll}\nu_{1}^{1}&\cdots&\nu_{1}^{p}\\ \vdots&&\vdots\\ \nu_{n}^{1}&\cdots&\nu_{n}^{p}\end{array}\right)=\left(\begin{array}[]{lll}X_{1}(\nu^{1})&\cdots&X_{1}(\nu^{p})\\ \vdots&&\vdots\\ X_{n}(\nu^{1})&\cdots&X_{n}(\nu^{p})\end{array}\right).$
@@
On the other hand, in view of $Y,X_{i}]=\sum_{s}\alpha_{i}^{s}X_{s}+\beta_{i}Y$ and the fact that $Y(\xi^{h})=Y(\nu^{j})=0$, last identity reduces to